A potential of 10000 V is applied across an xray tube.find the ratio of de Broglie wavelength associated with incident electrons to the min wavelength associated with xrays 
 
a. 10 
b. 20
C. 1\10
D. 1\20
 
I know this is simple but we have a terrible teacher at least he taught this badly. So can you explain please

Grade:12th Pass

1 Answers

Charchit Tailong
8 years ago
Lambda1=h/M1v1
________________
Lambda2=h/M2v2

L1/L2=is just a ratio of v2/v1
So by using value of pottential difference given,we can calculate velocities of different electrons